OpenStudy (anonymous):

solve the equation. Check the solution. -2/x+4 = 4/x+3

OpenStudy (anonymous):

\[\frac{ -2 }{ x+4 }=\frac{ 4 }{ x+3}\]

OpenStudy (anonymous):

answers : A.\[-\frac{ 13}{ 6 }\]

OpenStudy (anonymous):

B. -11

OpenStudy (anonymous):

C.\[-\frac{ 8 }{ 3 }\]

OpenStudy (anonymous):

D. \[-\frac{ 11 }{ 3 }\]

OpenStudy (anonymous):

@wyattjohnson

OpenStudy (anonymous):

multiply both sides by x

OpenStudy (anonymous):

it doesnt give me x though .. So how do i find x ?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

I'm sorry, I read your equation the way you wrote it at first, not the way you did the second time as -2/(x+4) = 4/(x+3) Just cross-multiply this way: -2(x+3) = 4(x+4) so -2x - 6 = 4x + 16 -22 = 2 x x = -11